Laubitz Company begins operations on April 1. Information from Job cost sheets shows the following:

Manufacturing Costs Assigned

Job Number

April

May

June

Month Completed

10

$6,500

$4,400

May

11

$4,100

$3,900

$3,000

June

12

$1,200

April

13

$4,700

$3,400

June

14

$5,400

$3,600

Not Complete

Each Job was sold for 30% above its cost in the month following completion.

Hint: Determine the month sold for each job and then complete the question. Remember to add the total costs of the jobs in April, May and June.

Use the information provided to answer the following questions:

a) What is the value of Laubitz's work-in-process inventory on June 30th? $
b) What is the value of Laubitz's finished goods inventory on June 30th? $

c) What is the value of Laubitz's gross profit for the month of July only? $
Answer #1

a) Value of Work in process inventory on June 30 = 5400+3600 = 9000

b) value of finished goods inventory on June 30 = 4100+3900+3000+4700+3400 = 19100

c) Gross profit = 19100*30% = 5730